# Physics in Next-token Prediction

### v1

@@ Abstract
We discovered the underlying physics in Next-token Prediction (NTP). We identified the law of information conservation within NTP and proposed the First Law of Information Capacity (IC-1), demonstrating that the essence of intelligence emergence in auto-regressive models is fundamentally a process of information transfer. We also introduced Landauer’s Principle into NTP, formulating the Second Law of Information Capacity (IC-2), which establishes the relationship between auto-regressive model training and energy consumption. Additionally, we presented several corollaries, which hold practical significance for production practices. Finally, we validated the compatibility and complementarity of our findings with existing theories.
@@ 1 Introduction
  • We validated the compatibility and complementarity of our perspectives with existing theories.

### v2

@@ Abstract
We discovered the underlying physics in Next-token Prediction (NTP). We identified the law of information conservation within NTP and proposed the First Law of Information Capacity (IC-1), demonstrating that the essence of intelligence emergence in auto-regressive models is fundamentally a process of information transfer. We also introduced Landauer’s Principle into NTP, formulating the Second Law of Information Capacity (IC-2), which establishes the relationship between auto-regressive model training and energy consumption. Additionally, we presented several corollaries, which hold practical significance for production practices. Finally, we demonstrate the consistency between the Law of Information Capacity and the Scaling Law for Neural Language Models, the Knowledge Capacity Scaling Laws, and the Scaling Laws for Precision.
@@ 1 Introduction
  • we demonstrate the consistency between our findings and the Scaling Law for Neural Language Models (Kaplan et al., 2020), the Knowledge Capacity Scaling Laws (Allen-Zhu & Li, 2024), and the Scaling Laws for Precision (Kumar et al., 2024).
@@ 4.4 Identify the Energy Limits Required for Training Auto-regressive Models
4.5 The conditions for lossless quantization
Quantizing the weights of autoregressive models to achieve economic goals is a common practice. However, achieving lossless quantization requires certain conditions. Suppose the model weights before quantization are $b$-bit with information capacity $\eta$, and after quantization, they are $b^{\prime}$-bit with information capacity $\eta^{\prime}$. To achieve lossless quantization, the following conditions must be met Eq. 11:
$\eta b\leq\eta^{\prime}b^{\prime}.$ (11)
Since $\eta^{\prime}\leq 1$, we can derive the necessary condition for lossless quantization as Eq. 12:
$\eta b\leq b^{\prime}.$ (12)
@@ 5.2 Consistency with the Knowledge Capacity Scaling Laws
5.3 Consistency with the Scaling Laws for Precision
(Kumar et al., 2024) found that as the model is trained on more data, the degradation caused by post-training quantization increases, ultimately making additional pre-training data detrimental. This is, in fact, an inevitable consequence of the First Law of Information Capacity. As the model is trained on more data, the information capacity $\eta$ of the model will gradually increase. When condition 12 is violated, quantization of the model will damage the information learned by the model, leading to negative effects.